OpenStudy (akshay_budhkar):

let w=e^{(i*pie)/3} and a,b,c,x,y,z be non zero complex numbers such that a+b+c=x a+bw+cw^2=y a+bw^2+cw=z thn the value of ([x]^2+[y]^2+[z]^2)/([a]^2+[b]^2+[c]^2) is??? where [q] is mod q...

OpenStudy (akshay_budhkar):

question from iit 2011

OpenStudy (anonymous):

ohkay

OpenStudy (akshay_budhkar):

w is cube root of unity i believe.. the complex cube root of unity

OpenStudy (anonymous):

What's iit?

OpenStudy (anonymous):

yeah it is

OpenStudy (akshay_budhkar):

IIT-indian institute of technology

OpenStudy (anonymous):

it is the best institute in india for engineering

OpenStudy (anonymous):

Ok, thks, I didn't know...

OpenStudy (anonymous):

Not really an engineering question:-)

OpenStudy (anonymous):

yeah but after 12 th we have to give entrance test jee( i don't know the full form) it's from that

OpenStudy (akshay_budhkar):

i had solved the ques using a short cut to save time

OpenStudy (anonymous):

jee to get into iit

OpenStudy (akshay_budhkar):

jee-joint entrance exam

OpenStudy (anonymous):

Mind u, maybe u have to be an engineer to solve it..

OpenStudy (anonymous):

oh well let me try it well i get confused jee e for entrance or engineering lol

OpenStudy (akshay_budhkar):

just take a=b=c=1 and get x,y,z and ur done!!!!! lol!!!! but now i want the technical answer

OpenStudy (anonymous):

well 3a = x+y+z

OpenStudy (akshay_budhkar):

that did get me 4 marks in the paper..substitution

OpenStudy (akshay_budhkar):

right add all the three. 1+w+wsquare is 0

OpenStudy (akshay_budhkar):

any suggestions estudier??? ishaan i cant find a way ahead of it..

OpenStudy (anonymous):

me too but wait a minute

OpenStudy (anonymous):

I don't like this kind of problems..:-

OpenStudy (anonymous):

well |3a| = |x + y+z|

OpenStudy (akshay_budhkar):

dude do u remeber (a+b+c)(a+bw+cw^2)(a+bw^2+cw)=?????? its a formula.. i am not remembering

OpenStudy (anonymous):

but x^2 makes the possible minus vanish and hence |x^2 + y^2+x^2| = x^2 + y^2 + z^2 \ just wait like 20 sec

OpenStudy (anonymous):

a^2 + b^2 +c^2 -ab +ac -cb

OpenStudy (anonymous):

not including (a+b+c)

OpenStudy (akshay_budhkar):

-ac dude.. all signs were same no?

OpenStudy (anonymous):

well including a+b+c it becomes x^3 + y^3 +z^3 -3xyz x=a y=b z=c

OpenStudy (akshay_budhkar):

ya right.. that was d formula..

OpenStudy (anonymous):

lets do it x^2 + y^2 +z^2 you find y while i find x^2 and z^2

OpenStudy (akshay_budhkar):

ok.. but i believe it is not so lenthy dude.. we get jus 3 minutes for a problem.. there must be some trick

OpenStudy (anonymous):

your right

OpenStudy (anonymous):

hey (x + y+z) ^2 = x2 +y2 + z2 +2xy + 2xz+2yz

OpenStudy (anonymous):

x + y +z = 3a

OpenStudy (anonymous):

hey (x + y+z) ^2 = x2 +y2 + z2 +2xy + 2xz+2yz I saw that in your puzzle from yesterday...

OpenStudy (anonymous):

9a^2 = x^2 + y^2 +z^2 +2xy + 2zy +2xz

OpenStudy (anonymous):

yeah

OpenStudy (akshay_budhkar):

right

OpenStudy (anonymous):

so ^2 make minus sign vanish

OpenStudy (anonymous):

it implies x^2 = |x|^2

OpenStudy (akshay_budhkar):

ofcourse.. its given just to confuse mod is not needed

OpenStudy (anonymous):

-ve sign don't exist in squares

OpenStudy (anonymous):

letme solve a lil

OpenStudy (anonymous):

:D

OpenStudy (akshay_budhkar):

i believe iit didnt want us to use the substitution method.... coz i have not found a way out till now after 3 months i gae the exam

OpenStudy (akshay_budhkar):

@face give it a try

OpenStudy (anonymous):

damn iit never let someone enter easily

OpenStudy (akshay_budhkar):

but the substitution method is alays there!!!!! lol!! the only weak point of iit jee.. lol!!!

OpenStudy (anonymous):

well what was the options

OpenStudy (akshay_budhkar):

*always

OpenStudy (anonymous):

lol

OpenStudy (akshay_budhkar):

its an integer type ques 0 to 9 all options

OpenStudy (anonymous):

aahhhh scary

OpenStudy (akshay_budhkar):

generally we mark 2 if we dun kno... 40 percent of ques hav ans 2 lol!!!

OpenStudy (anonymous):

Ok , Here I go

OpenStudy (akshay_budhkar):

ya go go bye. lol js kidding

OpenStudy (anonymous):

Okay Here we Have \[x = a+b+c\] \[y =a+bw+cw^2\] \[z = a + bw^2+cw\] Now we have to find \[\large{\frac{|x|^2 + |y|^2 + |z|^2 }{a^2 + b^2 +c^2}}\] \[|z|^2 = z*(conj.z)\] \[conj.y = a+bw^2 +cw\] \[conj.z = a+bw+cw^2\] Okay So we have now \[\frac{(a+b+c)^2 + 2(a + bw+cw^2)(a+bw^2+cw)}{a^2+b^2+c^2}\] \[\frac{3(a^2+b^2+c^2) +2(ab+bc+ac) - 2(ab+bc+ac)}{a^2+b^2+c^2}\]\[3\] Beautiful Solution : )

OpenStudy (akshay_budhkar):

Beauty